40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 1 11:31:50 2011, josephb, Update, CDS, MC1 LR digital reading close to zero, readback ~0.7 volts 
    Reply  Sun Jun 26 22:38:20 2011, rana, Update, CDS, MC1 LR dead for > 1 month; now revived temporarily lrsen.pngbt.jpg
       Reply  Mon Jun 27 00:23:11 2011, rana, Update, CDS, ETMX SIDE problem etmx-side.png
       Reply  Fri Nov 18 01:12:14 2011, Suresh, Update, CDS, MC1 LR dead for > 1 month; now revived temporarily 
Message ID: 4776     Entry time: Wed Jun 1 11:31:50 2011     Reply to this: 4888
Author: josephb 
Type: Update 
Category: CDS 
Subject: MC1 LR digital reading close to zero, readback ~0.7 volts 

There appears to be a bad cable connection somewhere on the LR sensor path for the MC1 optic.

The channel C1:SUS-MC1_LRPDMon is reading back 0.664 volts, but the digital sensor channel, C1:SUS-MC1_LRSEN_INMON, is reading about -16.  This should be closer to +1000 or so.

We've temporarily turned off the LRSEN filter module output while this is being looked into.

I briefly went out and checked the cables around the whitening and AA boards for the suspension sensors, but even after wiggling and making sure everything was plugged in solidly.  There was one semi-loose connection, but it wasn't on the MC1 board, but I pushed it all the way in anyways.  The monitor point on the AA board looks correct for the LR channels, although ITMX LR struck me as being very low at about -0.05 Volts.

According to data viewer, the MC1 LR sensor channel went bad roughly two weeks ago, around 00:40 on 5/18 UTC, or 17:40 on 5/17 PDT.

 

UPDATE:

It appears the AA board (or possibly the SCSI cable connected to it) is the problem in the chain.

ELOG V3.1.3-